The combustion of octane, C8H18, proceeds according to the reaction

2C8H18(l)+25O2(g)⟶16CO2(g)+18H2O(l)

If 474 mol of octane combusts, what volume of carbon dioxide is produced at 38.0∘C and 0.995 atm?
1
Expert's answer
2020-04-06T11:57:20-0400

V1 - volume of CO2 produced at T1= 273 K and p1=1atm.

V2 - volume of CO2 produced at T2=38.0∘C=311 K  and p2=0.995 atm


V1= 474*22.4=10617.6 L

p1*V1/T1=p2*V2/T2;

V2=(p1*V1*T2)/(p2*T1)

V2=(1*10617.6 *311)/(0.995*273)=12156.3 L;


Answer: 12156.3 L